# Generating synthetic data for neural operators

### v1

@@ 4 Numerical Experiments using the Fourier Neural Operator
The architecture we use for numerical experiments is the Fourier Neural Operator (FNO) introduced in (10), which is able to learn mappings between function spaces of infinite dimensions. The advantage of FNO is that it aims to approximate an operator that learns to solve a family of PDEs by mapping known parameters to the solution of that PDE, instead of only approximating one instance of a PDE problem. Due to the nature of FNO, this enables us to use our synthetic data in order to approximate an entire class of problems at once. The novelty of FNO is that the kernel function, which is learned from the data, is parameterized directly in Fourier space, leveraging the Fast Fourier Transform when computing the kernel function. We train FNO using Adam optimizer on batches of size $100$, with learning rate of $0.001$, modes set to $12$ and of width $64$. We also use relative $L_{2}$ error to measure performance for both training and testing.
We focus on second order semi-linear elliptic PDE equations in divergence form defined on $\Omega:=(0,1)^{2}$, with zero boundary conditions, given by
$-\operatorname{div}(A(x)\cdot\nabla u)+c(u)&=f&&\text{in }\Omega\\[5.0pt] B(u)&=0&&\text{on }\partial{\Omega},\end{aligned}\right.$ (7)
where $A(x)\in\mathbb{R}^{2\times 2}$, $c(u)$ is some nonlinear function in $u$ and $B(u)$ is either $B(u)=u$ or $B(u)=\nabla u\cdot\nu$, where $\nu$ is the exterior unit normal vector to the boundary $\partial\Omega$ that correspond to zero Dirichlet or Neumann condition, respectively. Here we also assume that $A$ is uniformly elliptic and each $a_{ij}\in L^{\infty}(\Omega)$ with $i,j\in\{1,2\}$. Note that, as a special case, for $A=I$ and $c(u)\equiv 0$, (7) reduces to the Poisson equation.
@@ 4.1 The Poisson Equation
First, we consider the simple example of problem (7), the Poisson equation
@@
$G^{\dagger}:L^{2}(\Omega)$ $\longrightarrow H(\Omega)$
@@
We now generate data points of the form $(f,u)$ where $u$ is defined as in (6) depending on $B(u)$, and $f$ is computed by taking derivatives of $u$ so that (8) holds. This way, we can generate a lot of data. Notice that the Poisson equation can be easily solved when fixing $f$, however, here we would like to demonstrate the method of generating data through the proposed method. As we will see later, obtaining solutions to the more general problem (7) is much more difficult, and in order to obtain data for neural operators, until now, classical numerical solvers have been used.
For the Poisson equation, we generate data points as described above and let $M$ in (6) range between $1$ and $20$, so that we can get a variety of such functions and various oscillations. We perform experiments by training with $1,000$, $10,000$ and $100,000$ functional data points and testing with $100$ data points for the Poisson problem with Dirichlet and then with Neumann boundary conditions. We report the relative $L_{2}$ errors in the following table 1.
@@
Training points
1,000
10,000
100,000
1,000
10,000
100,000
Training loss
0.02114
0.00413
0.00109
0.01431
0.00359
0.00123
Testing loss
0.05429
0.00533
0.00117
0.03341
0.00518
0.00127
@@
Testing on $f$ beyond finite trigonometric sums. Notice that if $u$ is represented as a finite linear sum of sines and cosines, as in (6), then $f$ generated according to (8) also consists of a finite linear sum of sines or cosines depending on $B$. Restricting our attention to the Dirichlet case, let us generate $f$ so that it does not consist of sine or cosine functions. This is akin to out-of-distribution testing in the machine learning literature. We consider the following two example functions: $f(x,y)=x-y$, which is smooth, and $f(x,y)=|x-0.5||y-0.5|,$ which has corners and is not everywhere differentiable in $\Omega$. However, in each case, $f$ is in $L^{2}(\Omega)$, and approximation of $L^{2}$ functions by trigonometric functions is well studied, and error bounds are available (see (5)). So we expect to obtain approximate solutions to the Poisson equation (8) for any $f$ function that is in $L^{2}(\Omega)$.
Figure 2: Ground truth of the Poisson equation when $f(x,y)=x-y$ (left) and $f(x,y)=|x-0.5||y-0.5|$ (right)
@@
In Figure $3$, we summarize the predicted solutions using FNO, when trained with $1,000$, $10,000$ and $100,000$ synthetic data functions that consist of sine functions given by (6). The following demonstrates that the choice of functions $u$ defined as in (6) generalizes well.
Figure 3: Predicted solutions using FNO when trained with $1,000$, $10,000$ and $100,000$ functional data points
Note that FNO performs better when predicting a solution to the Poisson equation when the right-hand side is given by a smooth function and has a harder time when the right-hand side has corners and is not everywhere differentiable in $\Omega$.
4.2 Second order linear elliptic PDE
@@
In problem (7), take $c(u)=0$ and allow the matrix $A$ to be of any form, possibly depending on $(x,y)$. Then (7) becomes
@@
$-\operatorname{div}(A\cdot\nabla u)&=f&&\text{in }\Omega\\[5.0pt] B(u)&=0&&\text{on }\partial{\Omega},\end{aligned}\right.$ (9)
@@
$G^{\dagger}:L^{2}(\Omega)$ $\longrightarrow H^{1}(\Omega)$
@@
Training points
1,000
10,000
100,000
1,000
10,000
100,000
Training loss
0.03545
0.00853
0.00319
0.03383
0.00840
0.00152
Testing loss
0.06309
0.01013
0.00333
0.09046
0.01146
0.00147
@@
$G^{\dagger}:L^{2}(\Omega)\times L^{\infty}(\Omega)\times L^{\infty}(\Omega)$ $\longrightarrow H(\Omega)$
@@
where $m_{i}$’s uniformly distributed in $[0.1,5]$ and $u$ is generated according to (6) with $M\in\{1,2,\dots,10\}$. For each generated data point, we generate a matrix of the above form and a function $u$ according to (6), then plug them both in equation (9) to compute $f$. Finally, the input data forms a triple $(f,\alpha,\delta)$, while the target is to predict $u$. This way, FNO learns how to solve a family of functions satisfying (9). We summarize the relative $L_{2}$ errors in table 3 using FNO when training with $1,000$, $5,000$ and $10,000$ data points. As we can see below and as expected, the performance of the FNO with more degrees of freedom in the input data is worse compared to the case where the matrix $A$ is considered fixed and held constant across all the input data.
@@
Training points
1,000
5,000
10,000
1,000
5,000
10,000
Training loss
0.12266
0.07352
0.03134
0.14295
0.04639
0.01107
Testing loss
0.27257
0.10641
0.04885
0.25906
0.07611
0.05508
@@
4.3 Second order semi-linear elliptic PDE
As a final experiment, we consider again (7) with $A=I$ and $c(u)=e^{2u}$. In this case, the problem becomes
$-\Delta u+e^{2u}&=f&&\text{in }\Omega\\[5.0pt] B(u)&=0&&\text{on }\partial{\Omega},\end{aligned}\right.$ (11)
As before, we generate $u$ as specified in (6) and compute $f$ by plugging into (11). Notice that we now have a non-linear term in $u$. However, from the theory, we know that the space of solutions is sill $H(\Omega)$. Numerical experiments show that despite the non-linearity in that term, we see that FNO achieves low $L_{2}$ relative errors, as indicated in table 4. We summarize the relative $L_{2}$ errors of training and testing loss in table 4 when we train on $1,000$, $10,000$ and $100,000$ data points and test on $100$ data points.
Dirichlet Neumann
Training points
1,000
10,000
100,000
1,000
10,000
100,000
Training loss
0.02310
0.00977
0.00372
0.01139
0.00432
0.00198
Testing loss
0.03141
0.00994
0.00352
0.03141
0.00996
0.00351
Table 4: FNO performance on the problem (11) using (6) functions.
5 Conclusion and future work

### v3

@@ 4 Numerical Experiments using the Fourier Neural Operator
In this section, we perform several numerical experiments to demonstrate the capabilities of our method. Through some linear and non-linear PDE examples, we show how our method generalizes for $f\in L^{2}(\Omega)$, see subsections 4.1 and 4.2. Namely, after training FNO with our data, we pick an $f$ that is not of trigonometric form, use a numerical solver to obtain a solution $u$ (at high resolution for better accuracy, which is then down-sample to $85\times 85$) and then see how this ("exact") solution $u$ compares to the solution predicted by FNO when trained only with our data. In subsection 4.3 we compare our method to a more classical approach of first generating $f$, then solving numerically for $u$.
The architecture we use for numerical experiments is the Fourier Neural Operator (FNO) introduced in [11], which can learn mappings between function spaces of infinite-dimensions. The advantage of FNO is that it aims to approximate an operator that learns to solve a family of PDEs by mapping known parameters to the solution of that PDE, instead of only approximating one instance of a PDE problem. Due to the nature of FNO, this enables us to use our synthetic data in order to approximate an entire class of problems at once. The novelty of FNO is that the kernel function, which is learned from the data, is parameterized directly in Fourier space, leveraging the Fast Fourier Transform when computing the kernel function. We train FNO using Adam optimizer on batches of size $100$, with a learning rate of $0.001$, modes set to $12$, and of width $64$. We also use relative $L_{2}$ error to measure performance for both training and testing.
We focus on second-order semi-linear elliptic PDE equations in divergence form defined on $\Omega:=(0,1)^{2}$, with zero boundary conditions, given by
$-\operatorname{div}(A(x)\cdot\nabla u)+b\cdot\nabla u+cu+g(u)&=f&&\text{in }\Omega\\[5.0pt] B(u)&=0&&\text{on }\partial{\Omega},\end{aligned}\right.$ (7)
where $A(x)\in\mathbb{R}^{2\times 2}$, $b\in\mathbb{R}^{2}$, $c\in\mathbb{R}$ and $g(u)$ is some nonlinear function in $u$ and $B(u)$ is either $B(u)=u$ or $B(u)=\nabla u\cdot\nu$, where $\nu$ is the exterior unit normal vector to the boundary $\partial\Omega$ that correspond to zero Dirichlet or Neumann condition, respectively. Here we also assume that $A$ is uniformly elliptic and each $a_{ij}\in L^{\infty}(\Omega)$ with $i,j\in\{1,2\}$.
@@ 4.1 The Poisson Equation
We first consider a simple example of problem (7), the Poisson equation, by taking $A(x)=I$ (the $2\times 2$ identity matrix), $b=(0,0)$, $c=0$ and $g(u)\equiv 0$
@@
$G:L^{2}(\Omega)$ $\longrightarrow H(\Omega)$
@@
Notice that the Poisson equation can be easily solved when fixing $f$, however, here we would like to demonstrate our method of generating data on this easy problem first. Later we will consider more complicated examples.
We generate data points of the form $(f,u)$ where $u$ is defined as in (6), depending on $B(u)$, and $f$ is computed by taking derivatives of $u$ so that (8) holds. This way, we can generate a lot of data. We let $M$ and $K$ in (6) range between $1$ and $20$, so that we can get a variety of such functions and various oscillations. We perform experiments by training with $1000$, $10000$, and $100000$ functional data points and testing with $100$ data points for the Poisson problem with Dirichlet and then with Neumann boundary conditions. We report the relative $L_{2}$ errors in the following Table 1.
@@
Training points 1,000 10,000 100,000 1,000 10,000 100,000
Training loss 0.01359 0.00322 0.00078 0.00818 0.00232 0.00065
Testing loss 0.02266 0.00346 0.00072 0.01877 0.00298 0.00066
@@
Testing on $f$ beyond finite trigonometric sums. Notice that if $u$ is represented as a finite linear sum of sines and cosines, as in (6), then $f$ generated according to (8) also consists of a finite linear sum of sines or cosines depending on $B$. So it is important to test on $f$’s that are not sums of sines or cosines to demonstrate that our method of generating data generalizes well.
@@
Restricting our attention to the Dirichlet case, let us generate $f$ so that it does not consist of sine or cosine functions. This is akin to out-of-distribution testing in the machine learning literature. We consider the following two example functions: $f_{1}(x,y)=x-y$, which is smooth, and $f_{2}(x,y)=|x-0.5||y-0.5|$, which is a not everywhere differentiable function. However, in each case, $f_{1},f_{2}$ are in $L^{2}(\Omega)$, and approximation of $L^{2}$ functions by trigonometric functions is well studied, and error bounds are available (see [5]). So we expect to obtain approximate solutions to the Poisson equation (8) for any $f$ function that is in $L^{2}(\Omega)$.
@@
In Figures 1 and 2, we summarize the predicted solutions using FNO, when trained with $1,000$, $10,000$ and $100,000$ synthetic data functions that consist of sine functions given by (6). We also record the relative $L^{2}$ errors for each example. The following demonstrates that the choice of functions constructed as in (6) generalizes well.
Figure 1: Predicted solutions of the Poisson equation (8) with $f_{1}(x,y)=x-y$ as the right-hand side using FNO with $1,000$, $10,000$ and $100,000$ training data points. Their relative $L^{2}$ errors are $0.097$, $0.096$ and $0.094$, respectively, while their relative $l^{\infty}$ errors are $0.145,0.123$ and $0.118$, respectively.
Figure 2: Predicted solutions of the Poisson equation (8) with $f_{2}(x,y)=|x-0.5||y-0.5|$ as the right-hand side using FNO with $1,000$, $10,000$ and $100,000$ training data points. Their relative $L^{2}$ errors are $0.102$, $0.114$ and $0.108$, respectively, while their relative $l^{\infty}$ errors are $0.326,0.188$ and $0.131$, respectively.
Note that FNO performs better when predicting a solution to the Poisson equation when the right-hand side is given by a smooth function, and has a harder time when the right-hand side is not smooth in $\Omega$. This behavior is expected from classical Fourier analysis. Smooth functions can be well-approximated by trigonometric polynomials, with the error decaying like $O(1\!/n^{k})$ if the function is $C^{k}$, where $n$ denotes the truncation number and $k$ denotes the number of continuous derivatives. For non-smooth functions, although pointwise convergence still holds, uniform convergence may fail and the approximation error decays much more slowly, explaining the drop in FNO performance from Figure 1 to Figure 2.
4.2 Second-order semi-linear elliptic PDE
We take $A=I$, $b=(0,0)$, $c=0$ and $g(u)=u^{2}$ in (7), in which case the problem becomes
$-\Delta u+u^{2}&=f&&\text{in }\Omega\\[5.0pt] B(u)&=0&&\text{on }\partial{\Omega},\end{aligned}\right.$ (9)
In this problem we have a nonlinear term $g(u)=u^{2}$ added. It turns out that despite the nonlinear term, we get decent approximations of solutions when using our data generation with FNO.
As before, we generate $u$ as specified in (6) and compute $f$ by plugging into (9). From the theory, we know that the space of solutions is $H(\Omega)$. Numerical experiments show that despite the non-linearity in that term, FNO achieves low $L_{2}$ relative errors, as indicated in Table 2. We summarize the relative $L_{2}$ errors of training and testing loss in Table 2 when we train on $1,000$, $10,000$ and $100,000$ data points and test on $100$ data points.
Dirichlet Neumann
Training points 1,000 10,000 100,000 1,000 10,000 100,000
Training loss 0.01679 0.01763 0.00184 0.01017 0.00800 0.00237
Testing loss 0.03391 0.02693 0.00562 0.02992 0.01472 0.00295
Table 2: FNO performance on the problem (9) using (6) functions.
Testing on $f$ beyond finite trigonometric sums. When we generate the unknown $u$ to be of sums of sines or cosines, when plugging in equation (9), the computed f still consists of sines and cosines, but with some terms squared. As before, after only training FNO with such $u$’s, we are interested in seeing how well it generalizes when testing non-trigonometric $L^{2}(\Omega)$ functions. We demonstrate generalization through the following two examples: $f_{1}(x,y)=xy$ and $f_{2}(x,y)=(x-0.5)^{2}+(y-0.5)^{2}$.
Figure 3: Predicted solutions of the semi-linear equation (9) with $f_{1}(x,y)=xy$ as the right-hand side using FNO with $1,000$, $10,000$ and $100,000$ training data points. Their relative $L^{2}$ errors are $0.058$, $0.058$ and $0.059$, respectively, while their relative $l^{\infty}$ errors are $0.089,0.084$ and $0.083$, respectively.
Figure 4: Predicted solutions of the semi-linear equation (9) with $f_{2}(x,y)=(x-0.5)^{2}+(y-0.5)^{2}$ as the right-hand side using FNO with $1,000$, $10,000$ and $100,000$ training data points. Their relative $L^{2}$ errors are $0.107$, $0.115$ and $0.118$, respectively, while their relative $l^{\infty}$ errors are $0.326,0.113$ and $0.106$, respectively.
The error plateaus after a certain amount of training data and stops decreasing further, even though the predicted solution becomes smoother. We notice similar behavior across several right-hand sides in equation (9) that smooth, albeit not finite sums of sines or cosines.
4.3 Comparison to a more classical approach
In this subsection, we compare our backward generation method for producing training data with a more classical forward approach based on solving the PDE numerically for randomly generated right-hand sides. Consider the same example as above, equation (9). We compare our method of generating data to using classical numerical solvers instead. In the forward generation method, we first draw $f\in L^{2}(\Omega)$ and then use a numerical solver to obtain a solution for the corresponding $u$.
We highlight a few difficulties with this approach: it is computationally more expensive to solve the PDE for each new $f$; it introduces additional error into the training data due to reliance on approximate numerical solutions; and in problems involving nonlinearity, like in example (9), uniqueness of solutions is not guaranteed.
To the best of our effort to provide a fair comparison, we draw $f$’s from $L^{2}(\Omega)$ by considering truncated series of orthonormal eigenvectors of $L^{2}(\Omega)$ with random coefficients. Recall the first part of Theorem 1, where eigenfunctions of the Laplace-Dirichlet operator form an orthonormal basis for $L^{2}(\Omega)$. Since we aim to solve (9) for every $f\in L^{2}(\Omega)$, it is natural to generate $f$’s as truncated linear combinations of such basis elements. Specifically, for any $f\in L^{2}(\Omega)$, we can write
$f(x,y)=\sum_{i,j=1}^{\infty}\langle f,e_{i,j}\rangle e_{i,j},$
with $e_{i,j}$ defined as in (3) or (4).
Intuitively, we are comparing two strategies: (1) approximating the $H_{0}^{1}$ space by generating representative $u$’s and directly computing the corresponding $f$ (our backward method), versus (2) approximating the $L^{2}$ space by drawing representative $f$’s and solving the PDE numerically to obtain $u$ (the forward method), which introduces additional numerical errors.
Our experiments indicate that our backward method is not only faster but also provides better accuracy when tested on non-trigonometric $f$’s. For the experiments, both training and testing data have resolution $85\times 85$, and we maintain the same resolution when generating training data via truncated series of basis functions in the forward method. It is important to note that discretization introduces error into the training set, and reducing this error requires increasing the grid resolution, which significantly slows down data generation. By contrast, our backward method does not introduce discretization error because the solutions are computed symbolically and exactly.
We trained the Fourier Neural Operator (FNO) model with $10,000$ training samples generated by each method. Data generation using our backward method took approximately $12$ minutes, leveraging the SymEngine library for symbolic differentiation. In contrast, data generation using the forward method required about $20$ minutes, relying on the FiPy Python library to numerically solve each instance on grids of size $85\times 85$.
After training on each dataset separately, we generate a set of $10$ functions $f$ used for testing. In order to obtain solutions as precise as possible, we solve for each $f$ on a $1700\times 1700$ grid and then downsample the results to $85\times 85$. The $f$’s are given by:
  1. 0.
    $2x(1-x)+2y(1-y)$
  2. 1.
    $xy$
  3. 2.
    $1$
  4. 3.
    $(x-0.5)^{2}+(y-1)^{2}$
  5. 4.
    Step function: $1$ if $x>0.5$, 00 otherwise
  6. 5.
    $x-y$
  7. 6.
    $|x-0.5||y-0.5|$
  8. 7.
    $(x-0.5)^{2}-(y-1)^{2}$
  9. 8.
    $\frac{1}{1+x^{2}+y^{2}}$
  10. 9.
    $e^{-\sqrt{x^{2}+y^{2}}}$
We summarize our findings in Figure 5.
Figure 5: Comparison of our backward generation method with the classical forward generation method, as described in Subsection 4.3, on the semi-linear problem (9). Relative $L^{2}$ errors of predicted solutions are shown: blue corresponds to our backward method and orange corresponds to the forward method. Our method consistently achieves lower errors across all test cases.
4.4 Second-order linear elliptic PDE
@@
In problem (7), take $g(u)=0$ and allow the matrix $A$ and the lower-order terms to be of any form, possibly depending on $(x,y)$. Then (7) becomes
@@
$-\operatorname{div}(A\cdot\nabla u)+b\cdot\nabla u+cu&=f&&\text{in }\Omega\\[5.0pt] B(u)&=0&&\text{on }\partial{\Omega},\end{aligned}\right.$ (10)
@@
$G:L^{2}(\Omega)$ $\longrightarrow H^{1}(\Omega)$
@@
Training points 1,000 10,000 100,000 1,000 10,000 100,000
Training loss 0.01992 0.01147 0.00262 0.03452 0.00621 0.00229
Testing loss 0.04780 0.01523 0.00274 0.08926 0.00848 0.00215
@@
$G:L^{2}(\Omega)\times L^{\infty}(\Omega)\times L^{\infty}(\Omega)$ $\longrightarrow H(\Omega)$
@@
where $m_{i}$’s are uniformly distributed in $[0.1,5]$ and $u$ is generated according to (6) with $M\in\{1,2,\dots,10\}$. For each generated data point, we generate a matrix of the above form and a function $u$ according to (6), then plug them both in equation (10) to compute $f$. Finally, the input data forms a triple $(f,\alpha,\delta)$, while the target is to predict $u$. This way, FNO learns how to solve a family of functions satisfying (10). We summarize the relative $L_{2}$ errors in Table 4 using FNO when training with $1,000$, $5,000$ and $10,000$ data points. As we can see below and as expected, the performance of the FNO with more degrees of freedom in the input data is worse compared to the case where the matrix $A$ is considered fixed and held constant across all the input data.
@@
Training points 1,000 5,000 10,000 1,000 5,000 10,000
Training loss 0.12266 0.07352 0.03134 0.14295 0.04639 0.01107
Testing loss 0.27257 0.10641 0.04885 0.25906 0.07611 0.05508
@@
4.5 Further examples
Second-order linear elliptic. We show an example of a linear second-order where we also include lower-order terms. For example take $A=I$, $b=(3,4)$, $c=1$ and $g(u)=0$. Then (7) becomes
$-\Delta u+3u_{x}+4u_{y}+u&=f&&\text{in }\Omega\\[5.0pt] B(u)&=0&&\text{on }\partial{\Omega},\end{aligned}\right.$ (12)
Once again, we would like to learn the operator
$G:L^{2}(\Omega)$ $\longrightarrow H^{1}(\Omega)$
$f$ $\longmapsto u$
Figure 7: Predicted solutions of the linear equation (12) with $f(x,y)=(x-0.5)^{2}-(y-0.5)^{2}$ as the right-hand side using FNO with 1,000, 10,000 and 100,000 training data points. Their relative $L^{2}$ errors are $0.287$, $0.156$ and $0.141$, respectively, while their relative $l^{\infty}$ errors are $0.440,0.251$ and $0.225$, respectively.
Additional semi-linear examples Here, we consider problem (7) with $A=I$, $b=(0,0)$, $c=0$, $g(u)=\varepsilon e^{u}$ and $B(u)=u$, that is
$-\Delta u+\varepsilon e^{u}&=f&&\text{in }\Omega\\[5.0pt] u&=0&&\text{on }\partial{\Omega},\end{aligned}\right.$ (13)
Notice that when $\varepsilon=0$, (13) becomes the Poisson equation. In this experiment, we let $\varepsilon\in[0,1]$ take values in increments of $0.1$ starting from 00. The purpose of this is to demonstrate the performance of our method as we go from a linear to a more non-linear problem by recalling the nonlinear term in (13).
For each $\varepsilon=0,0.1,\cdots,0.9,1.0$ we generate $10k$ training data and test on $100$, from which five examples are testing data where the right-hand side is first picked and we use a numerical solver to get the solution so we can test on whether we have good generalizations. We record the relative $L^{2}$ errors of the predicted solution to problem (13) by fixing the right-hand side $f$ and a $\varepsilon=0,0.1,\cdots,0.9,1.0$. We summarize the testing performance in the following plot and see that as we go from linear to non-linear, the performance improves, which at first seems surprising.
Figure 8: Plots of means of relative $L^{2}$ errors of 20 testing data for problem (13) generated using our method and 5 testing data with non-trig right-hand sides for which we invoked numerical solvers to obtain $u$. The five examples of non-trig $f$’s used on the above plot are given by: $x-y,xy,x^{2}+y-xy,2^{xy},20x-10y.$
The Poisson equation on a triangular domain. Eigenvalues and eigenfunctions of the Laplacian are also known in equilateral triangular domains given by $\Omega=\{(x,y)\in\mathbb{R}^{2}:0<x<1,0<y<\sqrt{3}x,y<\sqrt{3}(1-x)\}$, first discovered by Lamé [12] using reflection and symmetry arguments. We use the first $10$ eigenvalues (some are with multiplicity two) and their corresponding eigenfunctions in order to generate training data.
Figure 9: Predicted solutions of an example on a triangular domain using FNO with $1,000$ and $10,000$ training data. The relative $L^{2}$ errors are respectively $0.10233$ and $0.01424$.
5 Limitations, conclusion and future work
Limitations. Through experiments, we have observed that for certain “smooth" problems, our method generalizes well. However, as shown in the Darcy flow example (see Section 6.2), where the coefficients are non-smooth, our method did not generalize as effectively. While second-order elliptic PDEs represent a sizable class of problems, there are many other types of PDEs that fall outside this class, such as parabolic and hyperbolic. We have not yet tested the performance of our method on these other types of PDEs. We believe our method could apply to these other cases, but this requires further investigation. Finally, it is worth noting that the selected basis functions are not the only option, and different basis functions may be more suitable for certain problems. Alternative basis functions are worth investigating further.